Boat goes120 miles upstream in 20 hours and downstream in 10 hours. What is the speed of the boat and speed of the current?
Answer by stanbon(75887)

Boat goes120 miles upstream in 20 hours and downstream in 10 hours. What is the speed of the boat and speed of the current?
----
Upstream rate = 120/20 = 6 mph
---
Downstream rate = 120/10 = 12 mph
----
Equations:
Upstream rate: b - c = 6 mph
Downstream rate: b + c = 12 mph
-----
Add and solve for "b":
2b = 18 mph
boat speed = 9 mph
000
solve for "c":
b + c = 12
9 + c = 12
current speed = 3 mph
=========================
Cheers,
Stan H.
================